• Log InLog In
  • Register
Liquid`
Team Liquid Liquipedia
EST 03:26
CET 09:26
KST 17:26
  • Home
  • Forum
  • Calendar
  • Streams
  • Liquipedia
  • Features
  • Store
  • EPT
  • TL+
  • StarCraft 2
  • Brood War
  • Smash
  • Heroes
  • Counter-Strike
  • Overwatch
  • Liquibet
  • Fantasy StarCraft
  • TLPD
  • StarCraft 2
  • Brood War
  • Blogs
Forum Sidebar
Events/Features
News
Featured News
[ASL22] Finals Preview: Sweet Victory!12[ASL22] Ro4 Preview: Mirror Mirror6[ASL22] Ro8 Preview: Within Reach5[ASL22] Ro8 Preview: In A Tizzy11[ASL22] Ro16 Preview: Holy Diver5
Community News
Weekly Cups (Oct 26 - Nov 1): MaxPax, herO, Solar win0Weekly Cups (October 19-25): Clem triples first in 5.0.175Evil Geniuses - shutting down23ISSL (IntoTheTV x SOOP SC2 League): Premier7Weekly Cups (October 12-18): 5.0.17 era begins1
StarCraft 2
General
No more SC2 streamers on team liquid? Liquibets sc2 season 30 Evil Geniuses - shutting down Weekly Cups (Oct 26 - Nov 1): MaxPax, herO, Solar win How do you feel about the mass reverts in the 5.0.17 PTR?
Tourneys
Sparkling Tuna Cup - Weekly Open Tournament WardiTV Mondays ISSL (IntoTheTV x SOOP SC2 League): Premier Stellar Fest TWO the Moon (Dec 16-20) RSL goes to London! 2026 Offline Finals Nov 21-22
Strategy
[H] ZvP Mid-Late Game: Stalkers Collossi HT
Custom Maps
[D] the control group Simple Questions/Answers [G] How to playtest arcade maps [A] Nemrods 1/4 players
External Content
Mutation # 550 Graveyard Shift The PondCast: SC2 News & Results Mutation # 549 Endless Sparkles Mutation # 548 Shining Bright
Brood War
General
Rollback Netcode in Brood War BW General Discussion Top Players of October 2026 (EBPD Rankings) [Megathread] Daily Funny Videos Sagi.gg Launcher Released
Tourneys
KCM Race Survival 2026 Season 4 [Mixed-Gender CK] Rush+4 vs Best+4 New Worlds Star League [Megathread] Daily Proleagues
Strategy
50,000-Word Terran Builds Memo 2026-10-22 Cliff Jump Revisited (1 in a 1000 strategy) Simple Questions, Simple Answers
Other Games
General Games
Stormgate/Frost Giant Megathread Nintendo Switch Thread Deus Ex 3!? Mobile Legends: Bang Bang Armies of Exigo - YesYes?
Dota 2
Dota 2 Champions League Season 3 Begins April 25!
League of Legends
[TL LoL EUW IHs] Teemo shall perish
Heroes of the Storm
Heroes of the Storm 2.0
Hearthstone
Deck construction bug
TL Mafia
TL Mafia Community Thread
Community
General
US Politics Mega-thread Russo-Ukrainian War Thread Things Aren’t Peaceful in Palestine Canadian Politics Mega-thread The Chess Thread
Fan Clubs
Fan Club Resource Thread Reynor Fan Club The Flash Fan Club
Media & Entertainment
Anime Discussion Thread Movie Discussion! [Manga] One Piece
Sports
Football (Soccer) Thread TeamLiquid Health and Fitness Initiative For 2023 MLB/Baseball 2023
World Cup 2022
Tech Support
Computer Build, Upgrade & Buying Resource Thread
TL Community
Recent Gifted Posts
Blogs
Why I built Repalyzer (1.1.…
BlueStar
Innovate or Game Over?
TrAiDoS
ShieldBattery on Mac
Kraekkling
[ASL22] Ro4 Day1 Ticket Giv…
bITt.mAN
Can Bots Beat Pros?? Starcr…
namkraft
Customize Sidebar...

Website Feedback

Closed Threads



Active: 8285 users

The Law of Averages - Page 2

Blogs > obesechicken13
Post a Reply
Prev 1 2 All
MoonBear
Profile Blog Joined November 2010
Straight outta Johto18973 Posts
May 20 2011 22:57 GMT
#21
The statistical fallacies in this blog makes me sad. I understand how the message is supposed to work, but the execution was incredibly dubious.
ModeratorA dream. Do you have one that has cursed you like that? Or maybe... a wish?
Lemonwalrus
Profile Blog Joined August 2006
United States5465 Posts
May 20 2011 22:59 GMT
#22
Remake the blog, the more you try, the more likely you are to get it right sooner or later.

I learned that from a blog once.
EsX_Raptor
Profile Blog Joined February 2008
United States2802 Posts
May 21 2011 01:24 GMT
#23
Do you understand that there is a probability that you will never roll a 6 after an infinite amount of casts?
obesechicken13
Profile Blog Joined July 2008
United States10467 Posts
May 21 2011 02:15 GMT
#24
On May 21 2011 07:59 Lemonwalrus wrote:
Remake the blog, the more you try, the more likely you are to get it right sooner or later.

I learned that from a blog once.


Eh, you guys figured out what I was saying eventually.

On May 21 2011 10:24 EsX_Raptor wrote:
Do you understand that there is a probability that you will never roll a 6 after an infinite amount of casts?


No there isn't . pfft infinity.
I think in our modern age technology has evolved to become more addictive. The things that don't give us pleasure aren't used as much. Work was never meant to be fun, but doing it makes us happier in the long run.
Sleight
Profile Blog Joined May 2009
2471 Posts
Last Edited: 2011-05-21 02:25:36
May 21 2011 02:21 GMT
#25
Guys, you really don't understand probability theory.

Two dice, any chance of getting one single 6 is the same as rolling one die twice to get a six.

6 * 6 = 36 Outcomes
Winning outcomes = [6 X] and [X 6] with [6 6] only counting once, obviously, so 11 wins.
11 / 36 = .306

You guys need to do better math. Rolling more dice DOES increase the probability of a success. Taking more chances increases the likelihood of a success as a strategy.

EDIT for clarification: No single roll EVER has a higher likelihood than 1/6, but you can only examine probability on the large scale. We are looking at combinations of chances, which mean that if you roll 1000 dice, your likelihood of getting a 6 is basically 1, but that says nothing of whether its the first roll or the last roll. Probability theory ONLY applies to SETS.
One Love
Empyrean
Profile Blog Joined September 2004
17083 Posts
Last Edited: 2011-05-21 03:09:07
May 21 2011 03:03 GMT
#26
On May 21 2011 10:24 EsX_Raptor wrote:
Do you understand that there is a probability that you will never roll a 6 after an infinite amount of casts?


The probability is zero, because you will almost surely roll a 6 in an infinite amount of casts.

The amount of statistical ignorance in the general population is staggering

EDIT: Before someone replies, just because an event occurs with probability zero doesn't mean it can't occur. And I'm using the term "almost surely" in its technical meaning, in case there were any doubts.
Moderator
MoonBear
Profile Blog Joined November 2010
Straight outta Johto18973 Posts
May 21 2011 03:15 GMT
#27
On May 21 2011 11:21 Sleight wrote:
Guys, you really don't understand probability theory.

Two dice, any chance of getting one single 6 is the same as rolling one die twice to get a six.

6 * 6 = 36 Outcomes
Winning outcomes = [6 X] and [X 6] with [6 6] only counting once, obviously, so 11 wins.
11 / 36 = .306

You guys need to do better math. Rolling more dice DOES increase the probability of a success. Taking more chances increases the likelihood of a success as a strategy.

EDIT for clarification: No single roll EVER has a higher likelihood than 1/6, but you can only examine probability on the large scale. We are looking at combinations of chances, which mean that if you roll 1000 dice, your likelihood of getting a 6 is basically 1, but that says nothing of whether its the first roll or the last roll. Probability theory ONLY applies to SETS.

I'm not sure what point you are trying to make with your first example of rolling two dice concurrently. But if you are saying that the chance getting a 6 when rolling two dice at the same time is different from the probability of obtaining a 6 when rolling a single die is different then the result is a given. In that instance you are comparing a joint probability distribution to a single variable probability distribution so the results will be skewed as a result.

You are misunderstanding the point that Infinity was making. There is nothing wrong with observing probability on a small scale. Statisticians do this all the time.

The "Law of Large Numbers" (I do detest the name but no matter) is a statement that a very large number of observations will tend to reflect the underlying probability of the random variable in question. On this instance, I believe we are all in agreement. If you flip a fair coin an infinite number of times, your observations will tend to a 50:50 split.

However, you are confusing this with the applications and caveats that come with independence and random sampling. The OP is not making a statement about deducing conclusion from an extremely large sample space. Instead, it is a case of repeated trials and commenting on the results of each observation. In this case, it is a conditional probability function.

For example, one example was continuing to hit on women at a party until you achieve success. It would be incorrect to state that we are making observations on a large scale. Instead, you are taking repeated samples from a state space and then commenting on each observation one at a time. The fallacy in his article was "Given that I have made advances on a woman and failed, I expect to have a success because the law of averages should even things out eventually." This is incorrect because independence between events means that the conditional probability of success given failure is still the same as the probability of success in a single event.

I would recommend you read the link that Infinity posted and not make claims such as "You don't understand probability theory".
ModeratorA dream. Do you have one that has cursed you like that? Or maybe... a wish?
Sleight
Profile Blog Joined May 2009
2471 Posts
May 21 2011 03:25 GMT
#28
I pointed out exactly how the Gambler's fallacy doesn't account for reality, friend. I also said probability only works on SETS as my last statement. Two instances is a set, even if its largely useless. Two hundred instances and we begin to find relevant information to inform a decision, typically.

Notice how I explicitly say your single instance probably NEVER CHANGES?

Did you read my post before saying the same thing in a different way and trying to sound contradictory? I said nothing about the OP's correctness in regards to his Gambler's Fallacy. I spoke only to the event that increased trials increases chance of success. He is right, statistically you will eventually hit it home, but he is wrong in that it is some magical self-correction. It's just probability, nothing more, nothing less.

Like I said and you said and other informed people have said.
One Love
MoonBear
Profile Blog Joined November 2010
Straight outta Johto18973 Posts
May 21 2011 03:39 GMT
#29
On May 21 2011 12:25 Sleight wrote:
I pointed out exactly how the Gambler's fallacy doesn't account for reality, friend. I also said probability only works on SETS as my last statement. Two instances is a set, even if its largely useless. Two hundred instances and we begin to find relevant information to inform a decision, typically.

Notice how I explicitly say your single instance probably NEVER CHANGES?

Did you read my post before saying the same thing in a different way and trying to sound contradictory? I said nothing about the OP's correctness in regards to his Gambler's Fallacy. I spoke only to the event that increased trials increases chance of success. He is right, statistically you will eventually hit it home, but he is wrong in that it is some magical self-correction. It's just probability, nothing more, nothing less.

Like I said and you said and other informed people have said.

There are two disagreements with your argument I have.

The first is "Rolling more dice DOES increase the probability of a success." If we consider rolling die to be a Bernoulli or Binomial process, then each roll has the same probability. The probability of success on each roll hasn't changed. Increased trials does not "increase the chance of success" as you say. It simply reduces the effect of randomness and chance from your samples in the long run. What repeated samples means is that the distribution of your sample results will converge to the underlying distribution.

The second disagreement is that "you can only examine probability on the large scale". You can model any event regardless of how large your sample size is.
ModeratorA dream. Do you have one that has cursed you like that? Or maybe... a wish?
Sleight
Profile Blog Joined May 2009
2471 Posts
May 21 2011 04:03 GMT
#30
You have quit reading what I said, clearly.

Again, for the record. Increasing trials DOES NOT increase the probability of ANY SINGLE INSTANCE.

Got it? I've said that each and every post. Rolling a die 10k times does not increase the chance of getting a 6 on any given trial.

You are wrong to say that the probability of getting a 6 rolling a dice once is the same as rolling a dice 10k times. Dead wrong. Which is what I have said over and over. You are COMPLETELY WRONG if defining this as a Bernoulli process formally changes ANYTHING from what I've said.

Let's define our process as such. If a die rolls a non-6, we score it a 0. If the die is a 6 it is a 1, p = 1/6. Let's run a trial on that for one time.

We will say P(1) = (1 choose 1)*.16^1*.84*0 = .16. The chance of getting a 6 in 1 roll.

We will ask what are the chances of getting at least 1 6 in 2 trials.
P(1)=(2 choose 1)*.16^1*.84^1 = .268.

Does this account for all outcomes? NO! We still need to consider 2 successes.
P(2)=(2 choose2)*.16^2*.84^1 = .022

P of at least 1 success = .268+.022 = .29 (i rounded .166 to .16, so its marginally lower).

You are unequivocally wrong that changing the model somehow stops more trials from increasing chance of success.
One Love
obesechicken13
Profile Blog Joined July 2008
United States10467 Posts
Last Edited: 2011-05-21 04:37:51
May 21 2011 04:30 GMT
#31
On May 21 2011 12:15 MoonBear wrote:
Show nested quote +
On May 21 2011 11:21 Sleight wrote:
Guys, you really don't understand probability theory.

Two dice, any chance of getting one single 6 is the same as rolling one die twice to get a six.

6 * 6 = 36 Outcomes
Winning outcomes = [6 X] and [X 6] with [6 6] only counting once, obviously, so 11 wins.
11 / 36 = .306

You guys need to do better math. Rolling more dice DOES increase the probability of a success. Taking more chances increases the likelihood of a success as a strategy.

EDIT for clarification: No single roll EVER has a higher likelihood than 1/6, but you can only examine probability on the large scale. We are looking at combinations of chances, which mean that if you roll 1000 dice, your likelihood of getting a 6 is basically 1, but that says nothing of whether its the first roll or the last roll. Probability theory ONLY applies to SETS.

I'm not sure what point you are trying to make with your first example of rolling two dice concurrently. But if you are saying that the chance getting a 6 when rolling two dice at the same time is different from the probability of obtaining a 6 when rolling a single die is different then the result is a given. In that instance you are comparing a joint probability distribution to a single variable probability distribution so the results will be skewed as a result.

You are misunderstanding the point that Infinity was making. There is nothing wrong with observing probability on a small scale. Statisticians do this all the time.

The "Law of Large Numbers" (I do detest the name but no matter) is a statement that a very large number of observations will tend to reflect the underlying probability of the random variable in question. On this instance, I believe we are all in agreement. If you flip a fair coin an infinite number of times, your observations will tend to a 50:50 split.

However, you are confusing this with the applications and caveats that come with independence and random sampling. The OP is not making a statement about deducing conclusion from an extremely large sample space. Instead, it is a case of repeated trials and commenting on the results of each observation. In this case, it is a conditional probability function.

For example, one example was continuing to hit on women at a party until you achieve success. It would be incorrect to state that we are making observations on a large scale. Instead, you are taking repeated samples from a state space and then commenting on each observation one at a time.


Hmmm... ok... ok


The fallacy in his article was "Given that I have made advances on a woman and failed, I expect to have a success because the law of averages should even things out eventually." This is incorrect because independence between events means that the conditional probability of success given failure is still the same as the probability of success in a single event.

I would recommend you read the link that Infinity posted and not make claims such as "You don't understand probability theory".

Hey I never said that!
+ Show Spoiler +
;_; Even if it is true


And I quote
The law of averages is actually not a law at all. It's a truism that has no truth stating that when you've flipped a coin and gotten heads three times in a row, well then tails is more likely to turn up on your next flip. I just think that "The Law of Large Numbers" is a horrible title.


My intention wasn't to say that I would have a greater chance of success on my next attempt, I was saying that if I got turned down 20 times, and then approached 50 more women, I would have a higher chance of success with at least one of those 50 women than if I felt discouraged and only approached two women. I'm trying to say don't give up. I understand "independece between events" or whatever
+ Show Spoiler +
[image loading]
Hmpf

I think in our modern age technology has evolved to become more addictive. The things that don't give us pleasure aren't used as much. Work was never meant to be fun, but doing it makes us happier in the long run.
Cambium
Profile Blog Joined June 2004
United States16368 Posts
May 21 2011 04:41 GMT
#32
I was just going to type up something, then I saw two pages of comments

<3 TL
When you want something, all the universe conspires in helping you to achieve it.
infinity21 *
Profile Blog Joined October 2006
Canada6683 Posts
May 21 2011 05:44 GMT
#33
You guys are literally arguing over the difference between 'chances of success' and 'chances of a success' lol
Official Entusman #21
Empyrean
Profile Blog Joined September 2004
17083 Posts
May 21 2011 06:37 GMT
#34
On May 21 2011 14:44 infinity21 wrote:
You guys are literally arguing over the difference between 'chances of success' and 'chances of a success' lol


It's not worth trying.
Moderator
xxpack09
Profile Blog Joined September 2010
United States2160 Posts
May 21 2011 06:55 GMT
#35
You guys are missing the whole point of the blog. To use the dice analogy, the point is if you roll the die 100 times you're more likely to get a 6 in AT LEAST one of those 100 rolls than if you roll the die once. Which he applies to things like hooking up with girls, getting a job, etc.
Empyrean
Profile Blog Joined September 2004
17083 Posts
May 21 2011 09:56 GMT
#36
Right. But that's not what the law of averages is.
Moderator
Mactator
Profile Joined March 2011
109 Posts
Last Edited: 2011-05-21 10:35:07
May 21 2011 10:22 GMT
#37
On May 21 2011 13:30 obesechicken13 wrote:
Show nested quote +
On May 21 2011 12:15 MoonBear wrote:
On May 21 2011 11:21 Sleight wrote:
Guys, you really don't understand probability theory.

Two dice, any chance of getting one single 6 is the same as rolling one die twice to get a six.

6 * 6 = 36 Outcomes
Winning outcomes = [6 X] and [X 6] with [6 6] only counting once, obviously, so 11 wins.
11 / 36 = .306

You guys need to do better math. Rolling more dice DOES increase the probability of a success. Taking more chances increases the likelihood of a success as a strategy.

EDIT for clarification: No single roll EVER has a higher likelihood than 1/6, but you can only examine probability on the large scale. We are looking at combinations of chances, which mean that if you roll 1000 dice, your likelihood of getting a 6 is basically 1, but that says nothing of whether its the first roll or the last roll. Probability theory ONLY applies to SETS.

I'm not sure what point you are trying to make with your first example of rolling two dice concurrently. But if you are saying that the chance getting a 6 when rolling two dice at the same time is different from the probability of obtaining a 6 when rolling a single die is different then the result is a given. In that instance you are comparing a joint probability distribution to a single variable probability distribution so the results will be skewed as a result.

You are misunderstanding the point that Infinity was making. There is nothing wrong with observing probability on a small scale. Statisticians do this all the time.

The "Law of Large Numbers" (I do detest the name but no matter) is a statement that a very large number of observations will tend to reflect the underlying probability of the random variable in question. On this instance, I believe we are all in agreement. If you flip a fair coin an infinite number of times, your observations will tend to a 50:50 split.

However, you are confusing this with the applications and caveats that come with independence and random sampling. The OP is not making a statement about deducing conclusion from an extremely large sample space. Instead, it is a case of repeated trials and commenting on the results of each observation. In this case, it is a conditional probability function.

For example, one example was continuing to hit on women at a party until you achieve success. It would be incorrect to state that we are making observations on a large scale. Instead, you are taking repeated samples from a state space and then commenting on each observation one at a time.


Hmmm... ok... ok

Show nested quote +

The fallacy in his article was "Given that I have made advances on a woman and failed, I expect to have a success because the law of averages should even things out eventually." This is incorrect because independence between events means that the conditional probability of success given failure is still the same as the probability of success in a single event.

I would recommend you read the link that Infinity posted and not make claims such as "You don't understand probability theory".

Hey I never said that!
+ Show Spoiler +
;_; Even if it is true


And I quote
Show nested quote +
The law of averages is actually not a law at all. It's a truism that has no truth stating that when you've flipped a coin and gotten heads three times in a row, well then tails is more likely to turn up on your next flip. I just think that "The Law of Large Numbers" is a horrible title.


My intention wasn't to say that I would have a greater chance of success on my next attempt, I was saying that if I got turned down 20 times, and then approached 50 more women, I would have a higher chance of success with at least one of those 50 women than if I felt discouraged and only approached two women. I'm trying to say don't give up. I understand "independece between events" or whatever
+ Show Spoiler +
[image loading]
Hmpf



Oh I wish you were right and that success in life was proportional to the number of times you try. The problem though is that if you fuck up badly then the probability to succeed next time you try will decrease. Or the other way around. Real life events usually aren't independent and you can fail if you behave stupidly :-).

EDIT: One may say that this is true for sc2 as well. Trying a lot is a good thing but if you don't analyze your games, optimize your build orders and create new ones well.. then there is a chance that you aren't improving or perhaps even get bad habits.
Prev 1 2 All
Please log in or register to reply.
Live Events Refresh
Next event in 1h 34m
[ Submit Event ]
Live Streams
Refresh
StarCraft 2
Codebar 4
StarCraft: Brood War
Rain 4295
Shine 440
Noble 318
actioN 309
Killer 189
Hyuk 162
Free 149
firebathero 133
Leta 58
Mong 36
[ Show more ]
HiyA 35
ZergMaN 30
Sharp 18
ajuk12(nOOB) 7
yabsab 4
Dota 2
ODPixel354
monkeys_forever196
League of Legends
Reynor9
Other Games
summit1g10838
ceh9527
JimRising 492
Lowko215
Fuzer 159
Mew2King115
NeuroSwarm60
ViBE51
Railgan19
Organizations
Other Games
gamesdonequick565
Counter-Strike
PGL229
[ Show 12 non-featured ]
StarCraft 2
• AfreecaTV YouTube
• intothetv
• Kozan
• IndyKCrew
• Laughngamez YouTube
• Migwel
StarCraft: Brood War
• iopq 6
• BSLYoutube
• STPLYoutube
• ZZZeroYoutube
Dota 2
• lizZardDota2297
League of Legends
• Stunt464
Upcoming Events
The PondCast
1h 34m
KCM Race Survival
1h 34m
OSC
5h 34m
IntoTheTV X SOOP
2 days
Reynor vs Solar
ByuN vs SHIN
Clem vs MaxPax
herO vs Rogue
IPSL
2 days
WorsT vs nOmaD
Dewalt vs andersOn
AI Arena Tournament
2 days
IntoTheTV X SOOP
3 days
IPSL
3 days
Azhi_Dahaki vs ggnore13
Double vs WIZARD
Replay Cast
3 days
Replay Cast
4 days
[ Show More ]
Sparkling Tuna Cup
5 days
WardiTV Invitational
5 days
YoungYakov vs Lambo
SHIN vs TBD
Percival vs Shameless
Solar vs TBD
PiGosaur Cup
5 days
Liquipedia Results

Completed

Proleague 2026-11-04
Blizzard Classic Cup 2026
HCC Season 3

Ongoing

Acropolis #5
KCM Race Survival 2026 Season 4
International ShieldBattery League
BSL Season 23
IPSL Winter 2026-27
Calamity Cup Div. A Season 8
IEM Beijing 2026
PGL Masters Bucharest 2026
Stake Ranked Episode 5
1win Private Club #2
Thunderpick World Champ. '26
ESL Pro League Season 24
Stake Ranked Episode 4
1win Private Club #1
Logitech G Play Connect 2026
SL StarSeries Fall 2026
FISSURE Playground #3
BLAST Open Fall 2026
Esports World Cup 2026

Upcoming

1xBet Invitational
SC4ALL II: Brood War
BSL 23: Non-Korean Championship
HSC XXX
Stellar Fest 2: Lunar Cup
SC4ALL II: StarCraft II
Kung Fu Cup 2026 Grand Finals
RSL Offline Finals
Underdog Cup 2026 Div 2-3
PGL Bucharest 2027
BLAST Bounty Winter 2027
eXTREMESLAND 2026
PGL Major Singapore 2026
Stake Ranked Episode 6
BLAST Rivals Fall 2026
TLPD

1. ByuN
2. TY
3. Dark
4. Solar
5. Stats
6. Nerchio
7. sOs
8. soO
9. INnoVation
10. Elazer
1. Rain
2. Flash
3. EffOrt
4. Last
5. Bisu
6. Soulkey
7. Mini
8. Sharp
Sidebar Settings...

Advertising | Privacy Policy | Terms Of Use | Contact Us

Original banner artwork: Jim Warren
The contents of this webpage are copyright © 2026 TLnet. All Rights Reserved.